Can you please explain the following steps for solving the equation. I understand how to write it but need the steps for solving it.
One half of a number is 14 more than 2 times the number. Thank you

Answer by Theo(13342)

One half of a number is 14 more than 2 times the number.

let x = the number.

1/2 * x = 2*x + 14

subtract 2 * x from both sides of this eqaution to get:

1/2 * x - 2 * x = 14

2 * x is the same as 4/2 * x.

your equation can be written as 1/2 * x - 4/2 * x = 14

since 1/2 * x - 4/2 * x is equal to -3/2 * x, your equation becomes -3/2 * x = 14.

multiply both sides of this equation by -2/3 to get:

-2/3 * -3/2 * x = -2/3 * 14

since -2/3 * -3/2 = 1, your equation becomes:

x = -2/3 * 14

since -2/3 * 14 = -28/3, your equation becomes x = -28/3

that's your solution.

when x = -28/3, one half of a number is 14 more than 2 times the number becomes:

1/2 * (-28/3) = 2 * (-28/3) + 14

since 14 is the same as 42/3, this equation can be written as:

1/2 * (-28/3) = 2 * (-28/3) + (42/3)

this results in -14/3 = -56/3 + 42/3, which results in -14/3 = -14/3.

you get -14/3 = -14/3.

this confirms the solution is correct.

the solution is that x = -28/3.